OpenStudy (anonymous):

A side of one polygon is 18 meters and the corresponding side of a second similar polygon is 24 meters. If the perimeter of the first polygon is 60 meters and its area is 1000 square meters, find the perimter and area of the similar polygon.

OpenStudy (anonymous):

Primeter is proportional to (side) Area to (side)^2

OpenStudy (anonymous):

I don't get what you mean.

OpenStudy (anonymous):

primeter = (24/18)*60 area = (24/18)^2 * 1000

OpenStudy (anonymous):

does ^2 mean square root?

OpenStudy (anonymous):

NO square

OpenStudy (anonymous):

I mean squared!

OpenStudy (anonymous):

not square root, sorry.

jhonyy9 (jhonyy9):

18 --- 60 24 --- x x=(24*60)/18 = (24*10)/3 = 8*10 =80 will be the perimeter of secondly polygon area = 18 ---- 1000 24 ---- x x=(24*1000)/18 = (4*1000)/3 = 4000/3 =1333,3

OpenStudy (anonymous):

where did you get the 3 in the area?

jhonyy9 (jhonyy9):

simplify 24 and 18 by 6

OpenStudy (anonymous):

so you have to find the greatest common factor?

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

So the area for that other polygon is 1333.3?

jhonyy9 (jhonyy9):

after simplify end get the result of area

jhonyy9 (jhonyy9):

and

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

hope that is understandably

OpenStudy (anonymous):

So what is the formula you use? The other person's formula for finding the perimiter is similar to yours, but for the area, you guys have different formulas.

OpenStudy (anonymous):

I just don't understand how you use 3, if the gcf is 6. don't you divide by 6 instead of 3?

jhonyy9 (jhonyy9):

indifferent so if you see that for side 18 have got area equal 1000 so than for side 24 will be x

jhonyy9 (jhonyy9):

so 24 and 18 you can divide exactly by 6 24/6 =4 18/6 =3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Ohhhh yes! Got it.

OpenStudy (anonymous):

Thank you for your help jhonnyy!

jhonyy9 (jhonyy9):

my pleasure good luck bye